OpenStudy (anonymous):

jen is 13 years younger than andre . the sum of their ages in years is 137. what is andre's age in years? show work!

OpenStudy (amistre64):

j = a-13 j+a = 137 a-13+a = 137 2a = 150 a = 75

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

j = a-13 j+a = 137 a-13+a = 137 2a = 150 a = 75

OpenStudy (anonymous):

a medal.

OpenStudy (anonymous):

A medal for madbird2011.

OpenStudy (anonymous):

i would give you a medal but it looks like you copied amistre64. it the exact work amistre64 showed.